# Counterfactual Shapley Values for Explaining Reinforcement Learning

### v1

@@ Abstract
Introduction
Reinforcement learning (RL) has applications in autonomous navigation (?), healthcare (?), financial strategy optimization (?), and smart city management (?). However, the key to achieving widespread adoption of these applications lies in overcoming the challenge of explainability in RL models. the inability to adequately explain RL models could provoke ethical and regulatory issues and meet with resistance from potential users, collectively impeding the extensive application and further advancement of RL technology.
@@ Introduction
Shapley Value is a method for measuring the contribution of individuals to the total payoff in cooperative games. An application of Shapley values to Explainable Artificial Intelligence (XAI), each feature is regarded as a “player”, while the model’s predictive outcome is considered the “total payoff” (?). The Shapley Value assesses the average contribution of each feature to the predictive outcome by considering all possible combinations of features.
Investigations employing Shapley Values are remarkably limited across both multi-agent reinforcement learning (MARL) and single-agent in RL. However, MARL research demonstrates a relatively richer engagement with Shapley value applications, which are principally concentrated across three aspects. Firstly, Value Decomposition, (?) integrates Shapley Value theory with multi-agent Q-learning (SHAQ) to address value decomposition in global reward games, where Shapley Values play a critical role in ensuring equitable distribution of rewards based on individual agent contributions. Secondly, Credit Assignment, (?) introduces the Shapley Q-value within MARL (SQDDPG) to reallocate global rewards among agents, leveraging Shapley Values to quantify and fairly distribute the rewards reflecting each agent’s marginal contribution to the collective success. Lastly, Model Explanation, (?) the use of Shapley Values to explain cooperative strategies and individual contributions in multi-agent reinforcement learning, approximating Shapley Values via Monte Carlo sampling to reduce computational costs. Shapley Values offer a way to interpret complex multi-agent interactions, making it clearer how each agent’s decisions and actions contribute to the overall dynamics and results of the system.
To our knowledge, only the SVERL in (?), within the scope of RL research excluding MARL, utilizes Shapley Values to explain the decision-making process in reinforcement learning. Although SVERL provides a method for understanding and explaining the contribution of state features in the decision-making process of reinforcement learning agents, it has limitations in explaining specific action choices. The core of SVERL lies in analyzing the contribution of state features to the expected returns of an agent, rather than directly investigating why action $a$ is chosen over action $b$ in the same state, and it does not offer a mechanism to quantify the difference between these action choices. This means that while SVERL can help us indirectly understand how certain features influence an agent’s decisions, it cannot directly answer why an agent prefers a specific action, nor can it compare the merits of different actions. This limitation is not unique to SVERL but is a challenge faced by all methods that rely on Shapley Values to explain the decision-making process in reinforcement learning.
@@
In reinforcement learning, two main generative explanation methods by counterfactuals - one based on deep generative models and the other on generating counterfactual states - share a common limitation: the difficulty in quantifying differences between counterfactual instances. While methods based on deep generative models (?) can create realistic “what if” scenarios, comparing these instances in a high-dimensional latent space is challenging, as conventional distance metrics may not apply. On the other hand, methods that focus on minor state adjustments to guide different decisions also face challenges in quantifying the actual impact of these subtle changes on agent behavior, since even small modifications can have widespread effects in complex RL environments. Therefore, although these methods provide valuable insights into the decision-making process in RL, their difficulty in quantifying differences between counterfactual instances limits their application scope and the depth of their explanations.
To address the challenges of quantifying differences in reinforcement learning explanations, we develop a novel approach called “Counterfactual Shapley Value”. While previous research has explored the integration of counterfactual explanations with Shapley Values, such as addressing credit assignment problems in MARL in (?) or providing explanations in supervised learning (?), the former develops a novel method to assign credits in multi-agent systems using counterfactual thinking and Shapley values, which quantifies each agent’s contribution by considering what would happen if certain agents were absent, allowing for a fair and accurate assessment of each agent’s impact on the collective outcome, the latter uses counterfactual scenarios to enhance Shapley value explanations, providing insights on how changes to input features could influence model predictions. It makes explanations more actionable by showing which features to adjust to achieve desired outcomes. Our work is fundamentally different. For the first time, we have applied the combination of counterfactual reasoning and Shapley Values to the domain of reinforcement learning, introducing the innovative CSV approach. This method not only retains the benefits of traditional Shapley Values in quantitative analysis but also enhances the understanding of decision-making by examining specific hypothetical scenarios, such as “What would happen if a specific action were not taken?”. Our approach brings a more transparent and interpretable perspective to decision-making in reinforcement learning, significantly alleviating the limitations of previous explanatory methods.
@@
PRELIMINARIES
Markov Decision Process (MDP)
@@
Shapley Values in RL
In the domain of cooperative game theory, Shapley Values, originally proposed by (?), serve as a pivotal solution concept for the equitable allocation of collective gains or losses among participating entities, predicated on their individual contributions to the cumulative outcome. When transposed to the Reinforcement Learning (RL) framework, this concept facilitates an analytical assessment of the individual contributions of state features to the decision-making efficacy of an RL agent.
In RL, let $F=\{0,1,\ldots,n-1\}$ represent the set of indices corresponding to the comprehensive enumeration of state features within the state space $S$. The state space itself is structured into $n$ discrete dimensions, expressed as $S=S^{(0)}\times S^{(1)}\times\ldots\times S^{(n-1)}$, where each $S^{(i)}$ captures a distinct aspect of the environmental dynamics. Within this framework, each state is an ordered set, defined by $s=\{s^{(i)}|s^{(i)}\in S^{(i)},i\in F\}$. Let $C\subseteq F$ denote a subset of state features, specifically excluding the feature indexed by $i$. A partial observation of the state is then given by $s^{C}=\{s^{(i)}|s^{(i)}\in S^{(i)},i\in C\}$, which offers a perspective of the state based on a subset of observable features. The Shapley value (?) for a state feature $i$, denoted as $\phi_{i}$, quantifies the contribution of feature $i$ and is mathematically expressed as follows:
@@
Characteristic Value Function
@@
COUNTERFACTUAL SHAPLEY VALUES
@@
Average Counterfactual Difference
@@
Counterfactual Characteristic Value Function
@@
EXPERIMENTS
@@
Explanation in Gridworld 1 and Gridworld 2
@@
(c) $V(s)$, $Q(s,a)$ and $U(s)$ based on $\pi^{*}$ in GridWorld-1
(d) $V(s)$, $Q(s,a)$ and $U(s)$ based on $\pi^{*}$ in GridWorld-2
Figure 1: In a deterministic gridworld environment, the agent moves through a grid using actions: Up ($\uparrow$)↑ )), Down ($\downarrow$), Left ($\leftarrow$), and Right ($\rightarrow$). Each grid cell is a unique state, with the initial state marked in orange and the goal in green. Actions result in a -2 reward, with a +15 reward for reaching the goal (G). The discount factor $\gamma$ is 1. State features are the x and y coordinates, and red arrows indicate the optimal actions.
@@
Explanation in FrozenLake
FrozenLake is a reinforcement learning environment inspired by the challenge of traversing a frozen lake while avoiding treacherous holes, popularized by the OpenAI Gym framework (?). This environment is depicted as a grid where each cell is marked as either safe ice (Squares with Digits) or a perilous hole (H), with designated starting (S) and goal (G) positions. The agent’s objective is to navigate from the start to the goal without falling into any holes, guided solely by its position on the grid. The permissible actions are moving north, south, east, or west. The reward system is straightforward: falling into a hole incurs a -10 penalty, reaching the goal yields a +10 reward, and each move costs -1, incentivizing the agent to find the most efficient route to the goal. An episode begins with the agent at the start point and concludes successfully upon reaching the goal or unsuccessfully if the agent falls into a hole.
@@
Explanation in Taxi
The Taxi environment, a reinforcement learning benchmark by (?) and featured in OpenAI Gym (?), involves a taxi navigating a 5x5 grid to pick up and drop off a passenger at locations marked Red (R), Green (G), Blue (B), and Yellow (Y). The state captures the taxi’s position and the passenger’s status, either at a location or in the taxi. The taxi’s tasks are to pick up the passenger and deliver them to their destination. It has six actions: move in four directions, pick up, and drop off. Rewards are set to encourage efficiency: -1 for each move, +20 for successful delivery, and -10 for failed pick-up or drop-off, guiding the taxi towards effective passenger transit.
@@
Explanation in Pendulum
The Pendulum environment simulates an inverted pendulum swingup problem from (?), where the goal is to apply torque to swing the pendulum into an upright position. The system starts with a pendulum at a random angle ($\theta$) between $[-\pi,\pi]$ and a random angular velocity (torque) between $[-1,1]$. The observation space consists of the pendulum’s x-y coordinates at the free end and its angular velocity,state is $[cos(\theta),sin(\theta),torque]$, while the action space is a single-dimensional array between $[-2,2]$ representing the applied torque. This setup challenges control algorithms to stabilize the pendulum with minimal effort in a continuous state and action space.

### v2

# Explaining Reinforcement Learning: A Counterfactual Shapley Values Approach

@@ Abstract
Keywords:
Explainable Artificial Intelligence Explainable Reinforcement Learning Shapley Values.
1 Introduction
Reinforcement learning (RL) has applications in autonomous navigation [1], healthcare [2], financial strategy optimization [3], and smart city management [4]. However, the key to achieving widespread adoption of these applications lies in overcoming the challenge of explainability in RL models. the inability to adequately explain RL models could provoke ethical and regulatory issues and meet with resistance from potential users, collectively impeding the extensive application and further advancement of RL technology.
@@ 1 Introduction
Shapley Value is a method for measuring the contribution of individuals to the total payoff in cooperative games. An application of Shapley values to Explainable Artificial Intelligence (XAI), each feature is regarded as a “player”, while the model’s predictive outcome is considered the “total payoff” [5]. The Shapley Value assesses the average contribution of each feature to the predictive outcome by considering all possible combinations of features.
Investigations employing Shapley Values are remarkably limited across both multi-agent reinforcement learning (MARL) and single-agent in RL. However, MARL research demonstrates a relatively richer engagement with Shapley value applications, which are principally concentrated across three aspects. Firstly, Value Decomposition, [6] integrates Shapley Value theory with multi-agent Q-learning (SHAQ) to address value decomposition in global reward games, where Shapley Values play a critical role in ensuring equitable distribution of rewards based on individual agent contributions. Secondly, Credit Assignment, [7] introduces the Shapley Q-value within MARL (SQDDPG) to reallocate global rewards among agents, leveraging Shapley Values to quantify and fairly distribute the rewards reflecting each agent’s marginal contribution to the collective success. Lastly, Model Explanation, [8] the use of Shapley Values to explain cooperative strategies and individual contributions in multi-agent reinforcement learning, approximating Shapley Values via Monte Carlo sampling to reduce computational costs. Shapley Values offer a way to interpret complex multi-agent interactions, making it clearer how each agent’s decisions and actions contribute to the overall dynamics and results of the system.
To our knowledge, only the SVERL in [9], within the scope of RL research excluding MARL, utilizes Shapley Values to explain the decision-making process in reinforcement learning. Although SVERL provides a method for understanding and explaining the contribution of state features in the decision-making process of reinforcement learning agents, it has limitations in explaining specific action choices. The core of SVERL lies in analyzing the contribution of state features to the expected returns of an agent, rather than directly investigating why action $a$ is chosen over action $b$ in the same state, and it does not offer a mechanism to quantify the difference between these action choices. This means that while SVERL can help us indirectly understand how certain features influence an agent’s decisions, it cannot directly answer why an agent prefers a specific action, nor can it compare the merits of different actions. This limitation is not unique to SVERL but is a challenge faced by all methods that rely on Shapley Values to explain the decision-making process in reinforcement learning.
@@
In reinforcement learning, two main generative explanation methods by counterfactuals - one based on deep generative models and the other on generating counterfactual states - share a common limitation: the difficulty in quantifying differences between counterfactual instances. While methods based on deep generative models [10] can create realistic “what if” scenarios, comparing these instances in a high-dimensional latent space is challenging, as conventional distance metrics may not apply. On the other hand, methods that focus on minor state adjustments to guide different decisions also face challenges in quantifying the actual impact of these subtle changes on agent behavior, since even small modifications can have widespread effects in complex RL environments. Therefore, although these methods provide valuable insights into the decision-making process in RL, their difficulty in quantifying differences between counterfactual instances limits their application scope and the depth of their explanations.
To address the challenges of quantifying differences in reinforcement learning explanations, we develop a novel approach called “Counterfactual Shapley Value”. While previous research has explored the integration of counterfactual explanations with Shapley Values, such as addressing credit assignment problems in MARL in [11] or providing explanations in supervised learning [12], the former develops a novel method to assign credits in multi-agent systems using counterfactual thinking and Shapley values, which quantifies each agent’s contribution by considering what would happen if certain agents were absent, allowing for a fair and accurate assessment of each agent’s impact on the collective outcome, the latter uses counterfactual scenarios to enhance Shapley value explanations, providing insights on how changes to input features could influence model predictions. It makes explanations more actionable by showing which features to adjust to achieve desired outcomes. Our work is fundamentally different. For the first time, we have applied the combination of counterfactual reasoning and Shapley Values to the domain of reinforcement learning, introducing the innovative CSV approach. This method not only retains the benefits of traditional Shapley Values in quantitative analysis but also enhances the understanding of decision-making by examining specific hypothetical scenarios, such as “What would happen if a specific action were not taken?”. Our approach brings a more transparent and interpretable perspective to decision-making in reinforcement learning, significantly alleviating the limitations of previous explanatory methods.
@@
2 PRELIMINARIES
2.1 Markov Decision Process (MDP)
@@
2.2 Shapley Values in RL
In the domain of cooperative game theory, Shapley Values, originally proposed by [13], serve as a pivotal solution concept for the equitable allocation of collective gains or losses among participating entities, predicated on their individual contributions to the cumulative outcome. When transposed to the Reinforcement Learning (RL) framework, this concept facilitates an analytical assessment of the individual contributions of state features to the decision-making efficacy of an RL agent.
In RL, let $F=\{0,1,\ldots,n-1\}$ represent the set of indices corresponding to the comprehensive enumeration of state features within the state space $S$. The state space itself is structured into $n$ discrete dimensions, expressed as $S=S^{(0)}\times S^{(1)}\times\ldots\times S^{(n-1)}$, where each $S^{(i)}$ captures a distinct aspect of the environmental dynamics. Within this framework, each state is an ordered set, defined by $s=\{s^{(i)}|s^{(i)}\in S^{(i)},i\in F\}$. Let $C\subseteq F$ denote a subset of state features, specifically excluding the feature indexed by $i$. A partial observation of the state is then given by $s^{C}=\{s^{(i)}|s^{(i)}\in S^{(i)},i\in C\}$, which offers a perspective of the state based on a subset of observable features. The Shapley value [13] for a state feature $i$, denoted as $\phi_{i}$, quantifies the contribution of feature $i$ and is mathematically expressed as follows:
@@ 2.2 Shapley Values in RL
2.3 Characteristic Value Function
@@
3 COUNTERFACTUAL SHAPLEY VALUES
@@ 3.1 Counterfactual Differences
3.2 Average Counterfactual Difference
@@
3.3 Counterfactual Characteristic Value Function
@@
Figure 1: Computation of Counterfactual Shapley Value using VANI-CVF, CD-CVF, and ACD-CVF
An example of calculating CSV using vani-CVF, CD-CVF, and ACD-CVF, assuming a state with two features, is shown in Fig (1). In each CVF, each policy network is independently trained, and the resulting policy $\pi^{C}$ is learned based on the optimal policy for the current input dimensions $C$.
4 EXPERIMENTS
@@
4.1 Explanation in Gridworld 1 and Gridworld 2
@@
(c) $V(s)$, $Q(s,a)$, and $U(s)$ based on $\pi^{*}$ in GridWorld-1
(d) $V(s)$, $Q(s,a)$, and $U(s)$ based on $\pi^{*}$ in GridWorld-2
Figure 2: Comparison of GridWorld-1 and GridWorld-2, including their corresponding values of $V(s)$, $Q(s,a)$, and $U(s)$ based on $\pi^{*}$.
@@
4.2 Explanation in FrozenLake
FrozenLake is a reinforcement learning environment inspired by the challenge of traversing a frozen lake while avoiding treacherous holes, popularized by the OpenAI Gym framework [14]. This environment is depicted as a grid where each cell is marked as either safe ice (Squares with Digits) or a perilous hole (H), with designated starting (S) and goal (G) positions. The agent’s objective is to navigate from the start to the goal without falling into any holes, guided solely by its position on the grid. The permissible actions are moving north, south, east, or west. The reward system is straightforward: falling into a hole incurs a -10 penalty, reaching the goal yields a +10 reward, and each move costs -1, incentivizing the agent to find the most efficient route to the goal. An episode begins with the agent at the start point and concludes successfully upon reaching the goal or unsuccessfully if the agent falls into a hole.
@@ 4.2 Explanation in FrozenLake
4.3 Explanation in Taxi
The Taxi environment, a reinforcement learning benchmark by [15] and featured in OpenAI Gym [14], involves a taxi navigating a 5x5 grid to pick up and drop off a passenger at locations marked Red (R), Green (G), Blue (B), and Yellow (Y). The state captures the taxi’s position and the passenger’s status, either at a location or in the taxi. The taxi’s tasks are to pick up the passenger and deliver them to their destination. It has six actions: move in four directions, pick up, and drop off. Rewards are set to encourage efficiency: -1 for each move, +20 for successful delivery, and -10 for failed pick-up or drop-off, guiding the taxi towards effective passenger transit.
@@ 4.4 Explanation in Minesweeper
4.5 Explanation in Pendulum
The Pendulum environment simulates an inverted pendulum swingup problem from [16], where the goal is to apply torque to swing the pendulum into an upright position. The system starts with a pendulum at a random angle ($\theta$) between $[-\pi,\pi]$ and a random angular velocity (torque) between $[-1,1]$. The observation space consists of the pendulum’s x-y coordinates at the free end and its angular velocity,state is $[cos(\theta),sin(\theta),torque]$, while the action space is a single-dimensional array between $[-2,2]$ representing the applied torque. This setup challenges control algorithms to stabilize the pendulum with minimal effort in a continuous state and action space.